Find the slope of the line through ( -2,5) and (1,-7).
answer is : -4
need a full reak down of how the answer is -4 what has to be done to the integers in order for the answer to be -4

Answer by Mathtut(3670)

slope is rise over the run or the change in y over the change in x
:
(5-(-7))/(-2-1)=12/-3=-4 or (-7-5)/(1-(-2))=-12/3=-4
:
direction is important if you start with the first y coordinate for the change in y then you must also start with the first x coordinate for the change in x
and if you start with the 2nd y coorinate for the change in y then you MUST start with the 2nd x coordinate for the change in x.